OpenStudy (anonymous):

I'm stuck. I'm taking Geometry on FLVS and I'm studying for the final exam. I feel pretty confident about most of it except for the part where it asks me to "Complete the Square" to find the center and radius of a circle using a right triangle. So how do I factor an equation? Step 1) x^2-4x+y^2-2y=4-->Original Step 2) (x^2-4x)+(y^2-2y)=4-->Group like terms Step 3) (x^2-4x+4)+(y^2-2y+1)=4=(4+1)-->Complete quadratics Step 4) (x^2-4x+4)+(y^2-2y+1)=9-->Simplfiy Step 5) Answer: (x-2)^2+(y-1)^2=3^2-->Factor quadratics Step 6) Center-(2,1) Radius-3 This is mine, (x^2-4x+4)=(y^2-2y=1)=9

OpenStudy (anonymous):

@iambatman @Luigi0210 @Nnesha @bohotness

OpenStudy (anonymous):

@Mimi_x3, help plz:)

OpenStudy (mathstudent55):

Step 6 should read: Step 3) (x^2-4x+4)+(y^2-2y+1)=4+(4+1)-->Complete quadratics ^ plus instead of equal

OpenStudy (mathstudent55):

Your last line does not make sense. Why do you have all those equal signs? There should only be one equal sign. "This is mine, (x^2-4x+4)=(y^2-2y=1)=9"

OpenStudy (anonymous):

are those steps given ? so you do not understand how to get step 5 as you agree step 4 well look at step 5 and multiply it out you get your step 4 so it is now in circle equation form (x-h)^2 +(y-k)^2=r^2 (h,k) is the centre r^2=9 ....r=3